OpenStudy (anonymous):

factor completely 4x^2+12x+8

OpenStudy (ajprincess):

Wud u like me to explain the concept of factorising of these types of expressions?

OpenStudy (anonymous):

yes absolutely... i want to find out where i did these wrong, i figure them out on my end and then post it to see if i am right and then compare

OpenStudy (anonymous):

sum = 12 Product = 32

OpenStudy (anonymous):

find two numbers

OpenStudy (ajprincess):

let me explain it with a simple example so that u can get a gud idea of it. Suppose u hav an expression 2x^2+5x+3 and u can see that this is of the form ax^2+bx+c. Nw ur a=2. b=5 and c=3. 1.Multiply a by c and see vat u get. In this case u will get a*c=2*3=6 Say a*c=d=6 2.Find two numbers x and y such that x+y=b xy=d In this case x+y=5 xy=6 and x bcomes 3 and y bcomes 2. 3.Split the middle term. 2x^2+5x+3 =2x^2+2x+3x+3 4.let us take the first two terms and factorise them. 2x^2+2x as u can see 2x is common to both 2x^2 and 2x. When u take it out u get x+1. That is 2x^2+2x=2x(x+1) let us take the last two terms and factorise them. 3x+3 as u can see 3 is common to both 3x and 3. When u take it out u get x+1. That is 3x+3=3(x+1) 5.Nw let us combine both the results together. 2x^2+2x+3x+3 =2x(x+1)+3(x+1) =(2x+3)(x+1) 6.You r done. Is that clear? Hope u can do ur question with this idea in mind.

OpenStudy (anonymous):

Probably 8 and 4

OpenStudy (anonymous):

thanks yahoo but aj is helping me lol she is breaking it down for me :)

OpenStudy (anonymous):



OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

you guys are super helpful.... and @ajprincess i get confused cuz in the session, they taught us 3 ways to solve a problem and its all one jumbled mess in my head. thanks to you as well @Yahoo!

OpenStudy (ajprincess):

oh k. my kind advice is stick to one method nd it will bcome simple

OpenStudy (anonymous):

if only it were that simple... they gave us 3 methods and you have to figure out which method goes to each problem.. just too much

OpenStudy (ajprincess):

The one I explained is applicable for all questions of this type. If u think it simple nd feel it is really helpful u can stick to it.